# Pre-trained Transformer Uncovers Meaningful Patterns in Human Mobility Data

### v1

@@ Appendix A Appendix
A.1. Data processing
A.1.1. From GPS coordinates to trajectories
Initially, we projected the GPS data points onto an Uber H3 grid (Brodsky, 2018) of resolution 8 (Spatial grid of hexagon-shaped cells of an approximate area of 0.74 km2 each). Subsequently, sequences of data points are transformed into check-ins, defined as clusters of points belonging to the same user occurring within a 10-minute period and confined to the same cell. We constructed trajectories for each user by chronologically concatenating their cell hashes generated over a month’s time. To enhance data quality, we excluded trajectories with fewer than 10 check-ins or those with less than 3 unique grid cells. Additionally, for privacy considerations, we excluded any user identifying information, such user ID, from the final data.
A.1.2. Data splits
@@
A.2. Transformers
A.2.1. BERT
We adopted the Hugging Face (Wolf et al., 2019) implementation of BERT as the base transformer333See Table 14 for details on hyper-parameters.. Additionally, we experimented with two variations of BERT: 1) BERT initialized with random weights (BERTRandomly-initialized), and 2) BERT pre-trained on trajectory data (BERTPre-trained) in three different flavors: zero-shot, few-shots and fine-tuned.
@@ A.2.1. BERT
A.2.2. Baselines
We used the Hugging Face library to implement all three baselines (TransformerSmall, Medium & Large) detailed in Table 13. We limited our baselines to transformer based since transformers represent state-of-the-art in sequential data modeling. In all experiments we trained baselines in a fully supervised fashion using the exact same datasets utilized to fine-tune the BERT-based model.
@@ A.2.2. Baselines
A.2.3. Spatial tokenization
A.3. Pre-training and adaptation
A.3.1. Pre-training
First, whole-hash pre-masking with a 20% masking ratio is applied to both the training and validation data splits. Next, BERT was pre-trained, to minimize perplexity (Rabiner, 1989), for 40 epochs (Roughly 240 hours). Pre-training was done in parallel on two NVIDIA Tesla V100 GPU chips with 128G of memory each. See Table 14 for details on hyper-parameters.
A.3.2. Adaptation
To adapt a transformer-based model (BERT variants or baselines) we added on top a classification or regression layer (Depending on the task) initialized with random weights and subsequently trained the model on hand-labeled randomly-selected adaptation dataset for 20 epochs. The adaptation dataset consists of 50k samples444Depending on the target downstream task, a sample is either a hexagon randomly sampled from within Japan’s level-zero administrative boundaries or a trajectory randomly sampled from the test split. randomly split into 80% training and 20% testing subsets. The whole training split is used for fine-tuning, while a random 64 samples are used for few-shot learning. No training is conducted in zero-shot learning. The same machine used for pre-training was used for adaptation. See Table 14 for details on hyper-parameters.
@@ A.3.2. Adaptation
A.4. Evaluation
A.4.1. Evaluation metrics
To evaluate the quality of pre-training, we used perplexity (Rabiner, 1989) as a metric. Perplexity measures how ”surprised” the model is when it sees new data and is a commonly used metric in NLP in general and masked language modeling specifically.
@@ A.4.2. Pre-training
A.4.3. Adaptation
We reported the best results over 20 epochs while closely watching the validation and training loss curves for any signs of overfitting. Fine-tuning loss curves of all downstream tasks are shown in Figure 12.

### v2

@@ Appendix A Appendix
A-A Data processing
A-A1 From GPS coordinates to trajectories
Initially, we projected the GPS data points onto an Uber H3 grid [45] of resolution 8 (Spatial grid of hexagon-shaped cells of an approximate area of 0.74 km2 each). Subsequently, sequences of data points are transformed into check-ins, defined as clusters of points belonging to the same user occurring within a 10-minute period and confined to the same cell. We constructed trajectories for each user by chronologically concatenating their cell hashes generated over a month’s time. To enhance data quality, we excluded trajectories with fewer than 10 check-ins or those with less than 3 unique grid cells. Additionally, for privacy considerations, we excluded any user identifying information, such user ID, from the final data.
A-A2 Data splits
@@
A-B Transformers
A-B1 BERT
We adopted the Hugging Face [46] implementation of BERT as the base transformer333See Table XIV for details on hyper-parameters.. Additionally, we experimented with two variations of BERT: 1) BERT initialized with random weights (BERTRandomly-initialized), and 2) BERT pre-trained on trajectory data (BERTPre-trained) in three different flavors: zero-shot, few-shots and fine-tuned.
@@ A-B1 BERT
A-B2 Baselines
We used the Hugging Face library to implement all three baselines (TransformerSmall, Medium & Large) detailed in Table XIII. We limited our baselines to transformer based since transformers represent state-of-the-art in sequential data modeling. In all experiments we trained baselines in a fully supervised fashion using the exact same datasets utilized to fine-tune the BERT-based model.
@@ A-B2 Baselines
A-B3 Spatial tokenization
A-C Pre-training and adaptation
A-C1 Pre-training
First, whole-hash pre-masking with a 20% masking ratio is applied to both the training and validation data splits. Next, BERT was pre-trained, to minimize perplexity [48], for 40 epochs (Roughly 240 hours). Pre-training was done in parallel on two NVIDIA Tesla V100 GPU chips with 128G of memory each. See Table XIV for details on hyper-parameters.
A-C2 Adaptation
To adapt a transformer-based model (BERT variants or baselines) we added on top a classification or regression layer (Depending on the task) initialized with random weights and subsequently trained the model on hand-labeled randomly-selected adaptation dataset for 20 epochs. The adaptation dataset consists of 50k samples444Depending on the target downstream task, a sample is either a hexagon randomly sampled from within Japan’s level-zero administrative boundaries or a trajectory randomly sampled from the test split. randomly split into 80% training and 20% testing subsets. The whole training split is used for fine-tuning, while a random 64 samples are used for few-shot learning. No training is conducted in zero-shot learning. The same machine used for pre-training was used for adaptation. See Table XIV for details on hyper-parameters.
@@ A-C2 Adaptation
A-D Evaluation
A-D1 Evaluation metrics
To evaluate the quality of pre-training, we used perplexity [48] as a metric. Perplexity measures how ”surprised” the model is when it sees new data and is a commonly used metric in NLP in general and masked language modeling specifically.
@@ A-D2 Pre-training
A-D3 Adaptation
We reported the best results (Before overfitting) over 20 epochs while closely watching the loss curves for any signs of overfitting. Fine-tuning loss curves of all downstream tasks are shown in Figure 11.